OpenStudy (anonymous):

Help please??? Suppose you have a normally distributed set of data pertaining to a standardized test. The mean score is 1050 and the standard deviation is 200. What is the z-score of 1300 point score? 1.5 1.25 –0.25 –0.5

OpenStudy (kropot72):

The z-score is found using the following equation: \[\large z=\frac{X-\mu}{\sigma}=\frac{1300-1050}{200}=you\ can\ calculate\]

OpenStudy (anonymous):

Thank you <3

OpenStudy (kropot72):

You're welcome :)